How can you arrange the letters of the word SANDESH with vowels in alphabetical order 

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
Vowels are A and E
 
Hence it can be arranged in 7c2 *1 ways = 21 ways
 
And other letters can be arranged in 5! /2! Ways = 60 ways
 
Hence total ways to arrange ketters applying all conditions = 60 *21 = 1260